1. (12 points) In an arrangement for measuring the muzzle velocity of a rifle, the bullet is fired up at a wooden mass, into which it embeds (the bullet remains in the block). The wood is blasted straight up into the air to a measured height h. The mass of the bullet is 6.48 grams and the block is 4.81 kg. The block rose a height h-42 cm into the air from its initial position. a. (5 points) Determine the velocity of the block-bullet system just after the bullet embedded in the block b. (7 points) Determine the initial velocity of the bullet before it struck the block.
Answer #1

a) let the velocity of the Block- bullet system be v m/s

KE = PE

0.5*(M+m)v^2 = (M+m) gh

v = sqrt(2gh) = sqrt(2*9.8*0.042)

= 0.907 m/s

b) let the initial velocity of the bullet be v

by the conservation of momentum we have

mv = (M+m)V

0.00648*v = (4.81 + 0.00648)*0.907

v = 674.38 m/s